Please help. I am not good with fractions and adding this other stuff makes it even harder.
Complete the ordered pairs for the equation y = 2x + 4.

( ,0), (1/2, ), (0, ), (-1/2, )

Found 3 solutions by checkley75, tanimachatterjee, funmath:
Answer by checkley75(3666)

setting y=0 ( ,0) we get 0=2x+4, 2x=-4, x=-4/2, x=-2 thus we have (-2,0)
setting x=1/2 we get y=2*1/2+4, y=1+4, y=5 (1/2,5)
setting x=0 we get y=2*0+4, y=4 (0,4)
setting x=-1/2 we get y=2*-1/2+4, y=-1+4, y=3 (-1/2,3)

Answer by tanimachatterjee(60)

equation is y=2x+4............(1)
( , 0) that means y=0 given now put y=0 in the equn .(1)
we have 0=2x+4
or -2x=4
or -x= 4/2
or -x = 2
so x = -2
ans is (-2,0)
(1/2, ) that means x=1/2
so y= 2(1/2) +4
= 2/2 +4
= 1 +4
= 5
ans is ( 1/2 ,5)
(0, ) that means x=0
so y = 2*0 +4
or y = 4
ans is ( 0,4)
(-1/2 , ) here x= -1/2
so y = 2(-1/2) + 4
=-2/2 +4
= -1 + 4
=3
ans ( -1/2,3)

Answer by funmath(2933)

Complete the ordered pairs for the equation y = 2x + 4.

( ,0), (1/2, ), (0, ), (-1/2, )
Coordinates are in alphabetical order (x,y)
For ( ,0), let y=0 and solve for x.
0=2x+4
0-4=2x+4-4
-4=2x
-4/2=2x/2
-2=x
Therefore ( ,0)=(-2,0)
:
For (1/2, ), let x=1/2 and solve for y:
y=2(1/2)+4
y=2/2+4
y=1+4
y=5
Therefore (1/2, )=(1/2,5)
:
For (0, ), let x=0 and solve for y
y=2(0)+4
y=0+4
y=4
Therefore (0, )=(0,4)
:
For (-1/2, ) let x=-1/2 and solve for y
y=2(-1/2)+4
y=-2/2+4
y=-1+4
y=3
Therefore (-1/2, )=(-1/2,3)
Don't let the fractions scare you. Rule the fractions, don't let the fractions rule you. :-)
Happy Calculating!!!!